OpenStudy (anonymous):

The length of a rectangle is twice the width. If the length is increased by 4 inches and the width is diminished by 1 inch, a new rectangle is formed whose perimeter is 198 inches. Fine the dimensions of the original rectangle.

OpenStudy (anonymous):

width = x length = 2x If.... length 2x + 4 width x - 1 then L + w + L + w = 198 2x + 4 + x - 1 + 2x + 4 + x - 1 = 198 6x + 6 = 198 6x = 192 x = 32 width 32 length 64

OpenStudy (anonymous):

Length = 32 in Width = 16 in

OpenStudy (anonymous):

l=2w l+4 + 2(w+1)=198 l + 4 + 2w +2 = 198 -4 -4 l +2w +2 = 194 -2 -2 l +2w= 192 /2 /2 l +w= 96 l= 48 w= 24

OpenStudy (anonymous):

\[\{L=2w,2(L+4)+2(w-1)=198\} \]\[\{L=64,w=32\} \]